OpenStudy (anonymous):

Which equation is the perpendicular bisector of the line segment with endpoints (-2, 4) and (6, 8)? is it gonna be y=1/2x+5? i jus wanna make sure

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

y - 4 = -2(x+2) i guess

OpenStudy (anonymous):

because bisector, I would guess: y-6= -2*(x-2)...

OpenStudy (anonymous):

The perpendicular bisector would pass through the midpoint of the segment.

OpenStudy (anonymous):

y=-2x - 8